# What Ails Generative Structure-based Drug Design: Too Little or Too Much Expressivity?

### v1

@@ Abstract
Several generative models with elaborate training and sampling procedures have been proposed recently to accelerate structure-based drug design (SBDD); however, perplexingly, their empirical performance turns out to be suboptimal. We seek to better understand this phenomenon from both theoretical and empirical perspectives. Since most of these models apply graph neural networks (GNNs), one may suspect that they inherit the representational limitations of GNNs. We analyze this aspect, establishing the first such results for protein-ligand complexes. A plausible counterview may attribute the underperformance of these models to their excessive parameterizations, inducing expressivity at the expense of generalization. We also investigate this possibility with a simple metric-aware approach that learns an economical surrogate for affinity to infer an unlabelled molecular graph and optimizes for labels conditioned on this graph and molecular properties. The resulting model achieves state-of-the-art results using 100x fewer trainable parameters and affords up to 1000x speedup. Collectively, our findings underscore the need to reassess and redirect the existing paradigm and efforts for SBDD.
@@ 1 Introduction
Identifying new molecules or ligands that bind well to a protein target and have desired properties is a key challenge of Structure-based drug design (SBDD). Since experimental determination of binding is expensive and time-consuming, deep generative models offer hope to accelerate SBDD by rapidly suggesting good candidate molecules from a learned conditional distribution given the target protein, using experimental or in silico binding observations (Vamathevan et al., 2019; Bilodeau et al., 2022). Several approaches have been proposed for this purpose, including autoregressive models (Peng et al., 2022; Luo et al., 2021a), variational autoencoders (Ragoza et al., 2022), reinforcement learning (Li et al., 2021), and diffusion models (Schneuing et al., 2022; Guan et al., 2023a, b; Le et al., 2024).
@@
Despite a surge in interest from the machine learning community, these methods are observed to underperform empirically in terms of the docking scores, i.e., the estimated binding affinities of the candidates they generate. These suboptimal findings are particularly surprising since many of the aforesaid methods jointly optimize for atom types and 3D coordinates (often using GNN-based architectures for encoding), with intricate training and sampling procedures. Since the success of SBDD hinges on identifying candidates that dock well (Cieplinski et al., 2023), understanding this intriguing phenomenon holds key to utilizing the promise of generative modeling in SBDD.
@@
We take an initial but important step in this pursuit. Namely, we introduce and develop two contrasting hypotheses that shed light on different aspects of the problem. First, we investigate, theoretically, the possibility that the shortcomings of message-passing GNNs (Scarselli et al., 2009; Joshi et al., 2023), e.g., due to their inability to compute graph properties (Garg et al., 2020), might propagate to the protein-ligand complexes, yielding insufficient representations. It turns out that we can expose the limits on the expressivity of GNNs to distinguish ligands conditioned on the same protein target. To our knowledge, these are the first such results in conditional drug discovery contexts, motivating the design and analysis of more effective models that can mitigate the representational limitations.
Summary of our contributions
Theoretical limits of GNNs in SBDD (Section 3):
   GNNs limits extend to SBDD multi-body setting Proposition 2
   Real-world complexes differ in binding affinity but look identical to GNNs Appendix L
New model: SimpleSBDD (Section 4):
   Decomposition into unlabeled 2D molecular graph and atom labels Section 4.1
   Represent unlabelled graph with features unlearnable by GNNs Section 4.2
   Optimization of estimated binding built into model training Section 4.2
   Up to 1000x faster and 100x fewer parameters than baselines
Experiments (Section 5):
   Significantly higher estimated binding affinity than generative models Section 5.1
   Better tradeoff between estimated binding affinity and other
   molecular properties than optimization-based methods Section 5.2
   Drug repurposing: scoring 9000 molecules/second on CPU Section 5.2
Figure 1: Our contributions are theoretical, methodological and empirical.
@@
Second, we also study and analyze, empirically, a counterview that the generative models for SBDD might already be getting too sophisticated, owing, e.g., to their excessive parameterizations, and simpler methods may provide strong baselines. Some preliminary evidence on the efficacy of simple baselines for the unconditional molecule generation setting, but not SBDD, has emerged recently (Tripp and Hernández-Lobato, 2023) stemming from similar concerns.
@@
Concretely, we identify a key issue that has eluded attention heretofore; namely, that the existing methods predominantly focus on obtaining rich representations for drug properties such as QED with bulky models, while being oblivious to binding affinity as they seek to match the observed ligand distribution (which often includes molecules with suboptimal affinities). We address these issues with a novel simple two-phase generative method SimpleSBDD that draws inspiration from the success of performance-aware methods in other domains (Joachims, 2005), and optimizes for both the metrics of interest, i.e. molecular properties and estimated binding affinity.
@@
A key contribution of this work is topical: we draw attention to the disparity between the promise and the observed empirical performance of generative models for SBDD, outlining expressivity concerns. We also summarize our methodological, empirical and theoretical contributions in Figure 1.
@@ 2 Preliminaries and related work
Deep (Geometric) Learning for Drug Design. Deep learning has enabled progress on various facets of drug design across molecular generation (Jin et al., 2018; Zang and Wang, 2020; Verma et al., 2022; Shi et al., 2020; Luo et al., 2021b; Hoogeboom et al., 2022; Igashov et al., 2022), molecular optimization (Jin et al., 2019) and drug repurposing (Pushpakom et al., 2019). It has also accelerated advancements in protein folding and design (Jumper et al., 2021; Wu et al., 2022b; Ahdritz et al., 2022; Verkuil et al., 2022; Hie et al., 2022; Shi et al., 2023; Watson et al., 2022; Ingraham et al., 2022; Wu et al., 2022a; Verma et al., 2023) and docking (Stärk et al., 2022b; Ganea et al., 2022; Corso et al., 2023). We refer the reader to the surveys by Chen et al. (2018) and Pandey et al. (2022).
Graph Neural Networks (GNNs) have emerged as a workhorse for modeling data, such as molecules, that can be represented as graphs (Scarselli et al., 2009; Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018). They can be extended to encode important symmetries in geometric graphs (Satorras et al., 2021), so have found success across diverse tasks in the drug design pipeline including docking (Corso et al., 2023) and molecular property prediction (Stärk et al., 2022a).
However, GNNs have known representational limits (Xu et al., 2019; Morris et al., 2019; Maron et al., 2019; Garg et al., 2020; Wang and Zhang, 2022; Puny et al., 2023a; Joshi et al., 2023). There have been many methods proposed for improving them (Bouritsas et al., 2022; Vignac et al., 2020; Puny et al., 2023b) albeit potentially at the expense of generalization (Karczewski et al., 2024).
@@
The goal of SBDD is to generate drug candidates with high binding affinity to a target protein, while controlling for other properties. We list below common molecular properties, calculated in practice using the RDKit software (Landrum et al., 2020), that are relevant to drug design (Bilodeau et al., 2022; Peng et al., 2022): 1) Binding affinity – the strength of the connection between the ligand and the protein. A common surrogate for affinity is the Vina score (Alhossary et al., 2015); 2) LogP – logarithm of the partition coefficient, which is a solubility indicator; 3) QED – quantitative estimate of drug-likeness, a mixture of properties correlated with drugs; 4) SA – Synthetic accessibility, an estimate of how easy it would be to synthesize the molecule; and 5) Lipinski’s Rule of Five – A heuristic about whether a molecule would be an active oral drug.
@@
Setting. We view proteins as labelled graphs: each pocket $G=(V,E)$ consists of nodes $V=\{v_{1},\ldots,v_{N}\}$ with $v_{i}$ representing atoms, and edges $E\subseteq V\times V$. Each node $v=(a,\mathbf{s})$ is associated with an atom type $a\in\mathcal{A}=\{\texttt{C},\texttt{N},\texttt{O},\ldots\}$ and 3D coordinates $\mathbf{s}\in\mathbb{R}^{3}$, and each edge $e\in E$ with a bond type from $\{1,2,3\}$. We represent the molecule $M$ analogously. We include superscripts $G^{P}$ and $G^{M}$ to distinguish between protein and molecule graphs. We also distinguish between a labelled molecular graph $G^{M}$ and unlabelled one $U^{M}$. The unlabelled graph contains information about the number of nodes and edges between them, but no information about atom types or 3D coordinates.
@@ 3 Representation limits of GNNs for SBDD
The generative models for SBDD are quickly increasing in size (see Table 1) aiming to expand their expressivity. However, we bring attention to the representational issues concerning large models. Specifically, we show that there exist molecular graphs that cannot be distinguished by a large class of GNN models regardless of their depth, even in the additional protein context. Consequently, certain graph properties, including the number of cycles, diameter, and sizes of cycles, cannot be learned by message-passing graph neural networks (MPGNNs). Interestingly, as we will demonstrate in Section 4, these features are very useful in predicting binding affinity.
It is known that there exist graphs representing distinct molecular structures that cannot be distinguished using message-passing graph neural networks (MPGNNs) (Sato, 2020; Garg et al., 2020). Following the notation of Garg et al. (2020), we consider Locally Unordered GNNs (LU-GNNs), which we summarize as follows. The updated embedding $h_{v}^{(l)}$ of node $v$ at layer $l$ is defined as
$m_{u\to v}^{(l-1)}=\phi(h_{u}^{(l-1)},e_{uv});\tilde{h}_{v}^{(l-1)}=\mbox{AGG}\{m_{u\to v}^{(l-1)}\>|\>u\in N(v)\};h_{v}^{(l)}=\mbox{COMBINE}\{h_{v}^{(l-1)},\tilde{h}_{v}^{(l-1)}\},$
where $N(v)$ denotes the set of neighbours of $v$, $e_{uv}$ are edge features and $\phi$ is any function. It is known that there exist non-isomorphic graphs, for which LU-GNNs produce identical embeddings to every node at each layer (Garg et al., 2020). We now note that this result also holds for graphs embedded in 3D space. Let us introduce Locally Unordered 3D GNNs (LU3D-GNN) which aggregate embeddings for each node based on the embeddings of the neighbours, as well as their distance to these neighbours. Formally, in LU3D-GNNs, the messages are defined as:
@@ 4.1 Decoupling the unlabelled molecular graph from atom types
One of the main innovations underlying our framework is the separation of molecular representation into the unlabelled molecular graph and the labels, the atom types. Our motivation to adopt such an approach is primarily rooted in empirical findings that highlight the strong relationship between the structure of molecules and their predicted binding affinities, as approximated by the Vina software. In a preliminary experiment, we investigated how Vina binding scores behave when we modify the atom types and coordinates, while keeping the underlying graph intact.
Remarkably, our analysis showed a strong coefficient of determination, with an $R^{2}$ value of 0.76, between the scores of the original molecules and their modified counterparts. This significant result highlights that a considerable portion of the variability in predicted binding affinity can be attributed to the information contained within the unlabelled molecular graph itself, even before incorporating atom-specific details. Please see Appendix A for details. Note that many graph features characterizing the unlabeled molecular graph, such as the diameter and the number and sizes of rings, are unlearnable by LU-GNNs (Section 3). We will leverage this insight in defining our model.
It is important to recognize that these findings pertain to the Vina software and not the intrinsic biological process of binding. While Vina is the most prevalent approximation, it remains just that – an approximation. Nonetheless, upon conducting similar experiments with Gnina (McNutt et al., 2021), a reportedly significantly more accurate binding approximation than Vina, we found analogous results. This suggests that the observed behavior is not solely an artifact of Vina but may be more widespread. Please see Appendix A.4 for details.
Inspired by these findings, we propose a new model SimpleSBDD (see Figure 3). We first generate the high-binding unlabelled graph structures $U^{M}$, and generate atom types $\mathbf{a}^{M}$ independently of the protein. The model is defined to approximate the conditional distribution
@@
where we assume atoms $\mathbf{a}^{M}$ are conditionally independent of the protein pocket $G^{P}$ given the unlabelled graph structure $U^{M}$. This approximation does incur information loss as one would expect. However, as we will see in the next section, performance-aware learning mitigates this effect. The generative model thus reduces to two independent components: unlabelled graph sampler and atom sampler. We discuss them below.
@@ 4.2 Performance-aware learning
Typically, one would train the unlabelled graph model $p(U^{M}|G^{P})$ to match the empirical distribution. However, we hypothesize that observed ligand-protein complexes contain examples with suboptimal binding, and propose to focus on the most promising complexes with the following procedure. We first learn a binding surrogate $g_{\theta}$ to predict binding affinity from unlabelled graph alone,
@@
Next, we use $g_{\theta}$ to score unlabelled graphs from a repository of molecules, and choose those with the best predicted affinities. We define a generative distribution
$p_{\theta}(U^{M}|G^{P})=\mbox{UNIFORM}(\big{\{}U^{M}\in\mathcal{U}^{M}\>|\>g_{\theta}(U^{M},G^{P})\in\left[v_{\mathrm{min}},v_{\mathrm{max}}\right]\big{\}}),$ (5)
where $\mathcal{U}^{M}$ is a database of unlabelled graph structures, and $v_{\mathrm{min}},v_{\mathrm{max}}$ are threshold hyperparameters. We set $v_{\mathrm{min}}$ and $v_{\mathrm{max}}$ to be the 5th and 10th percentiles of all predictions for $\mathcal{U}^{M}$ to avoid outliers (see Appendix B for details). As $\mathcal{U}^{M}$, we used the ZINC250k (Irwin et al., 2012) instead of the CrossDocked2020 (Francoeur et al., 2020), which only contains 8K unique ligands.
We note that instead of sampling from a database, one can define a generative model over unlabelled graphs. We have not investigated this direction due to substantial variability in the molecular space even with predefined unlabelled graphs. Specifically, one can generate at least a thousand distinct and chemically diverse molecules sharing the same unlabelled graph (see Appendix C for details).
Architecture of the scoring model $g_{\theta}$
@@
To train the binding affinity predictor $g_{\theta}$, we need pairs of protein-ligand pairs with their affinity estimates. We chose not to use the CrossDocked2020 dataset, the gold standard for SBDD, for the following reasons: 1) even though it contains 100K protein-ligand pairs, there are only around 8K unique ligands, which can be insufficient for generalisation,
(a) Protein pocket with optimal ligand size $\approx 20$
(b) Protein pocket with optimal ligand size $\geq 38$
@@
We visualize the predictions of $g$ as a function of the size of the molecule. We notice that protein pockets fall under two cases: either there is an optimal size of the ligand (Figure 4(a)) or there is a monotonic relationship favouring larger ligands (Figure 4(b)). We hypothesize the model learnt the optimal ligand size per pocket, but for some pockets it is larger than the largest available molecule (38 atoms).

### v2

# What Ails Generative Structure-based Drug Design: Expressivity is Too Little or Too Much?

@@ Abstract
Several generative models with elaborate training and sampling procedures have been proposed to accelerate structure-based drug design (SBDD); however, their empirical performance turns out to be suboptimal. We seek to better understand this phenomenon from both theoretical and empirical perspectives. Since most of these models apply graph neural networks (GNNs), one may suspect that they inherit the representational limitations of GNNs. We analyze this aspect, establishing the first such results for protein-ligand complexes. A plausible counterview may attribute the underperformance of these models to their excessive parameterizations, inducing expressivity at the expense of generalization. We investigate this possibility with a simple metric-aware approach that learns an economical surrogate for affinity to infer an unlabelled molecular graph and optimizes for labels conditioned on this graph and molecular properties. The resulting model achieves state-of-the-art results using 100x fewer trainable parameters and affords up to 1000x speedup. Collectively, our findings underscore the need to reassess and redirect the existing paradigm and efforts for SBDD. Code is available at https://github.com/rafalkarczewski/SimpleSBDD.
@@ 1 INTRODUCTION
Identifying new molecules or ligands that bind well to a protein target and have desired properties is a key challenge of Structure-based drug design (SBDD). Since experimental determination of binding is expensive and time-consuming, deep generative models offer hope to accelerate SBDD by rapidly suggesting good candidate molecules for a given target protein, using experimental or in silico binding observations (Vamathevan et al.,, 2019; Bilodeau et al.,, 2022). Several approaches have been proposed, including autoregressive models (Peng et al.,, 2022; Luo et al., 2021a, ), variational autoencoders (Ragoza et al.,, 2022), reinforcement learning (Li et al.,, 2021), and diffusion models (Schneuing et al.,, 2022; Guan et al., 2023a, ; Guan et al., 2023b, ; Le et al.,, 2024).
@@
Despite a surge in interest from the machine learning community, these methods are observed to underperform empirically in terms of the docking scores, i.e., the estimated binding affinities of the candidates they generate. These findings are particularly surprising since many of the aforesaid methods jointly optimize for atom types and 3D coordinates (often using GNN-based architectures), with intricate training and sampling procedures. Since the success of SBDD hinges on identifying candidates that dock well (Cieplinski et al.,, 2023), understanding this phenomenon holds key to utilizing the promise of generative modeling in SBDD.
@@
We take an initial but important step in this pursuit by proposing two contrasting hypotheses that shed light on different aspects of the problem. First, we investigate, theoretically, the possibility that the shortcomings of message-passing GNNs (Scarselli et al.,, 2009; Joshi et al.,, 2023), e.g., due to their inability to compute graph properties (Garg et al.,, 2020), might propagate to the protein-ligand complexes, yielding insufficient representations. It turns out that we can expose the limits on the expressivity of GNNs to distinguish ligands conditioned on the same protein target. To our knowledge, these are the first such results in conditional drug discovery contexts, motivating the design and analysis of more effective models that can mitigate the representational limitations.
@@
Second, we also study and analyze, empirically, a counterview that the generative models for SBDD might already be getting too sophisticated, owing, e.g., to their excessive parameterizations, and simpler methods may provide strong baselines. Some preliminary evidence on the efficacy of simple baselines for the unconditional molecule generation setting, but not SBDD, has emerged recently (Tripp and Hernández-Lobato,, 2023) stemming from similar concerns.
@@
Concretely, we identify a key issue that has eluded attention heretofore; namely, that the existing methods predominantly focus on obtaining rich representations for drug properties such as QED with bulky models, while being oblivious to binding affinity as they seek to match the observed ligand distribution (which often includes molecules with suboptimal affinities). We address these issues with a novel simple two-phase generative method SimpleSBDD that draws inspiration from the success of performance-aware methods in other domains (Joachims,, 2005), and optimizes for both the metrics of interest, i.e. molecular properties and estimated binding affinity.
@@
1.1 Contributions of this work
We summarize our contributions below. This work
  1. 1.
    (Topical) draws attention to the disparity between the promise and the observed empirical performance of generative models for SBDD, outlining concerns pertaining to their expressivity;
  2. 2.
    (Theoretical) initiates a formal analysis for limits on the expressivity of methods for SBDD, exposing the representational challenges inherent in GNN models for encoding protein-ligand complexes;
  3. 3.
    (Methodological) introduces performance-aware optimization in SBDD settings, using a simple generative model that optimizes for both docking scores and physicochemical properties; and
  4. 4.
    (Empirical) demonstrates the versatility of the proposed approach SimpleSBDD using three instantiations tailored to problems of drug repurposing (suggesting candidates from a known ligand set), generation of novel ligands, and property optimization that seeks to control both molecular properties and binding. SimpleSBDD outperforms existing methods with orders of magnitude fewer parameters and faster runtime.
@@ 2 PRELIMINARIES AND RELATED WORK
Deep (Geometric) Learning for Drug Design. Deep learning has enabled progress on various facets of drug design across molecular generation (Jin et al.,, 2018; Zang and Wang,, 2020; Verma et al.,, 2022; Shi et al.,, 2020; Luo et al., 2021b, ; Hoogeboom et al.,, 2022; Igashov et al.,, 2022), molecular optimization (Jin et al.,, 2019) and drug repurposing (Pushpakom et al.,, 2019). It has also accelerated advancements in protein folding and design (Jumper et al.,, 2021; Wu et al., 2022b, ; Ahdritz et al.,, 2022; Verkuil et al.,, 2022; Hie et al.,, 2022; Shi et al.,, 2023; Watson et al.,, 2022; Ingraham et al.,, 2022; Wu et al., 2022a, ; Verma et al.,, 2023) and docking (Stärk et al., 2022b, ; Ganea et al.,, 2022; Corso et al.,, 2023). We refer the reader to the surveys by Chen et al., (2018) and Pandey et al., (2022).
Graph Neural Networks (GNNs) have emerged as a workhorse for modeling graph data, such as molecules (Scarselli et al.,, 2009; Kipf and Welling,, 2017; Hamilton et al.,, 2017; Veličković et al.,, 2018). They can be extended to encode important symmetries in geometric graphs (Satorras et al.,, 2021), so have found success across diverse tasks in the drug design pipeline including docking (Corso et al.,, 2023) and molecular property prediction (Stärk et al., 2022a, ).
However, GNNs have known representational limits (Xu et al.,, 2019; Morris et al.,, 2019; Maron et al.,, 2019; Garg et al.,, 2020; Wang and Zhang,, 2022; Puny et al., 2023a, ; Joshi et al.,, 2023). There have been improvements proposed (Bouritsas et al.,, 2022; Vignac et al.,, 2020; Puny et al., 2023b, ) albeit potentially at the expense of generalization (Karczewski et al.,, 2024).
@@
The goal of SBDD is to generate drug candidates with high binding affinity to a target protein, while controlling for other properties. We list below common molecular properties, calculated in practice using the RDKit software (Landrum et al.,, 2020), that are relevant to drug design (Bilodeau et al.,, 2022; Peng et al.,, 2022): 1) Binding affinity – the strength of the connection between the ligand and the protein. A common surrogate for affinity is the Vina score (Alhossary et al.,, 2015); 2) LogP – logarithm of the partition coefficient, which is a solubility indicator; 3) QED – quantitative estimate of drug-likeness, a mixture of properties correlated with drugs; 4) SA – Synthetic accessibility, an estimate of how easy it would be to synthesize the molecule; and 5) Lipinski’s Rule of Five – A heuristic about whether a molecule would be an active oral drug.
@@
Setting. We view proteins as labelled graphs: each pocket $G=(V,E)$ consists of nodes $V=\{v_{1},\ldots,v_{N}\}$ with $v_{i}$ representing atoms, and edges $E\subseteq V\times V$. Each node $v=(a,\mathbf{s})$ is associated with an atom type $a\in\mathcal{A}=\{\texttt{C},\texttt{N},\texttt{O},\ldots\}$ and 3D coordinates $\mathbf{s}\in\mathbb{R}^{3}$. We include superscripts $G^{P}$ and $G^{M}$ to distinguish between protein and molecule graphs. We also distinguish between a labelled molecular graph $G^{M}$ and unlabelled one $U^{M}$. The unlabelled graph contains information about the number of nodes and edges between them, but no information about atom types or 3D coordinates.
@@ 3 REPRESENTATION LIMITS OF GNNS FOR SBDD
The generative models for SBDD are quickly increasing in size (see Table 1) aiming to expand their expressivity. However, we bring attention to the representational issues concerning large models. Specifically, we show that there exist molecular graphs that cannot be distinguished by GNN models regardless of their depth, even in the additional protein context. Consequently, certain graph properties, including the number of cycles, diameter, and sizes of cycles, cannot be learned by message-passing graph neural networks (MPGNNs). Interestingly, as we will demonstrate in Section 4, these features are very useful in predicting binding affinity.
It is known that there exist distinct molecular structures that cannot be distinguished using message-passing graph neural networks (MPGNNs) (Sato,, 2020; Garg et al.,, 2020). Following the notation of Garg et al., (2020), we consider Locally Unordered GNNs (LU-GNNs), summarized as follows. The updated embedding $h_{v}^{(l)}$ of node $v$ at layer $l$ is defined as
$\begin{split}m_{u\to v}^{(l-1)}&=\phi(h_{u}^{(l-1)},e_{uv})\\ \tilde{h}_{v}^{(l-1)}&=\mbox{AGG}\{m_{u\to v}^{(l-1)}\>|\>u\in N(v)\}\\ h_{v}^{(l)}&=\mbox{COMBINE}\{h_{v}^{(l-1)},\tilde{h}_{v}^{(l-1)}\},\end{split}$
where $N(v)$ denotes the set of neighbours of $v$, $e_{uv}$ are edge features and $\phi$ is any function. It is known that there exist non-isomorphic graphs, for which LU-GNNs produce identical embeddings (Garg et al.,, 2020). We now note that this result also holds for graphs embedded in 3D space. Let us introduce Locally Unordered 3D GNNs (LU3D-GNN) which aggregate embeddings for each node based on the embeddings of the neighbours, as well as their distance to these neighbours. Formally, in LU3D-GNNs, the messages are defined as:
@@ 4.1 Decoupling the unlabelled molecular graph from atom types
One of the main innovations underlying our framework is the separation of molecular representation into the unlabelled molecular graph and the labels, the atom types. Our motivation to adopt such an approach is primarily rooted in empirical findings that highlight the strong relationship between the structure of molecules and their predicted binding affinities, as approximated by the Vina software. In a preliminary experiment, we investigated how Vina binding scores behave when we modify the atom types and coordinates while keeping the underlying graph intact.
Remarkably, our analysis showed a high correlation at $\rho=0.83$, between the scores of the original molecules and their modified counterparts. This result highlights that a considerable portion of the variability in predicted binding affinity can be attributed to the information contained within the unlabelled molecular graph itself, even before incorporating atom-specific details (details in Appendix A). Note that features characterizing the unlabeled molecular graph, such as the diameter and the number of rings, are unlearnable by LU-GNNs (Section 3). We will leverage this insight in our model.
It is important to recognize that these findings pertain to the Vina software and not the biological process of binding. While Vina is the most prevalent approximation, it remains just that – an approximation. Nonetheless, upon conducting similar experiments with Gnina (McNutt et al.,, 2021), a reportedly significantly more accurate binding approximation than Vina, we found similar results (Appendix A.3). This suggests that the observed behavior is not solely an artifact of Vina but may be more widespread.
Inspired by these findings, we propose a new model SimpleSBDD (Figure 2). We first generate the high-binding unlabelled graph structures $U^{M}$, and generate atom types $\mathbf{a}^{M}$ independently of the protein. The model is defined to approximate the conditional distribution
@@
where we assume atoms $\mathbf{a}^{M}$ are conditionally independent of the protein pocket $G^{P}$ given the unlabelled graph structure $U^{M}$. This approximation does incur information loss as expected. However, as we will see in the next section, performance-aware learning mitigates this effect. The generative model thus reduces to two independent components: unlabelled graph sampler and atom sampler. We discuss them below.
@@ 4.2 Performance-aware learning
Typically, the unlabelled graph model $p(U^{M}|G^{P})$ is trained to match the data distribution. However, we hypothesize that observed ligand-protein complexes contain examples with suboptimal binding, and propose to focus on the most promising complexes with the following procedure. We first learn a scoring model $g_{\theta}$ to predict binding affinity from unlabelled graph alone,
@@
To represent $U^{M}$ we use properties describing its structure such as its number of rings, number of rotatable bonds and graph diameter, which cannot be learned by MPGNNs (See Appendix D for details). Next, we use $g_{\theta}$ to score unlabelled graphs from a repository of molecules, and choose those with the best predicted affinities. We define a generative distribution
$\begin{split}&p_{\theta}(U^{M}|G^{P})=\mbox{UNIFORM}(U(G^{P})),\\ &U(G^{P})=\big{\{}U^{M}\in\mathcal{U}^{M}\>|\>g_{\theta}(U^{M},G^{P})\in\left[v_{\mathrm{min}},v_{\mathrm{max}}\right]\big{\}},\end{split}$ (5)
where $\mathcal{U}^{M}$ is a database of unlabelled graph structures, and $v_{\mathrm{min}},v_{\mathrm{max}}$ are threshold hyperparameters. We set $v_{\mathrm{min}}$ and $v_{\mathrm{max}}$ to be the 5th and 10th percentiles of all predictions for $\mathcal{U}^{M}$ to avoid outliers (Appendix B). As $\mathcal{U}^{M}$, we used the ZINC250k (Irwin et al.,, 2012) instead of the CrossDocked2020 (Francoeur et al.,, 2020), which only contains 8K unique ligands.
We note that instead of sampling from a database, one can define a generative model over unlabelled graphs. We have not investigated this due to substantial molecular variability even with predefined unlabelled graphs. Specifically, one can generate at least a thousand distinct and chemically diverse molecules sharing the same unlabelled graph (Appendix C).
@@
To train the scoring model $g_{\theta}$, we need pairs of protein-ligand pairs with their affinity estimates. We chose not to use the CrossDocked2020 dataset, the gold standard for SBDD, for the following reasons: 1) even though it contains 100K protein-ligand pairs, there are only around 8K unique ligands,
(a) Optimal ligand size $\approx 20$
(b) Optimal ligand size $\geq 38$
@@
We visualize the predictions of $g$ as a function of the size of the molecule. We notice that protein pockets fall under two cases: either there is an optimal size of the ligand (Figure 3(a)) or there is a monotonic relationship favouring larger ligands (Figure 3(b)). This is because the largest available molecule is 38 atoms, which is smaller than the learned optimum.